HUBBLE SPACE TELESCOPE DAILY REPORT #1764 PERIOD COVERED: 0000Z (UTC) 11/20/96 - 0000Z (UTC) 11/21/96 Daily Status Report as of 326/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.2 Completed FOS/BL/GHRS/1 6795 (Sleuthing the Dynamo: The Final = Frontier) The Faint Object Spectrograph (Blue Side) and the High Resolution = Spectrograph (Side One) were used to make ultraviolet observations of = late-type STAR-0330. This is to study the nature of its magnetic = activity, which is thought to be the link between the hydro-magnetic = dynamo deep inside the star and the violent state of its outer most = coronal layers. The observations were completed as planned, and no = problems were reported. 1.3 Completed WF/PC-2 7211 (WF/PC-2 Pre-SM97 UV Baseline Observation.) The WF/PC-2 was used to make far-ultraviolet observations of CVZ = Class target WD0214+568 as part of a pre-observation baseline of SM97 = (WF/PC-2 SMOV Calibration). The observations were completed as = scheduled, and no anomalies were noted. 1.4 Completed GHRS/1 6800 (Dueterium in the Interstellar Medium Toward = Hot Stars) The High Resolution Spectrograph (Side One) was used to make = spectrographic absorption observations of the interstellar medium (ISM) = along the line-of-sight toward the hot star HD48915. A wavelength = calibration observation was also taken. The deuterium abundance in the = local ISM toward Sirius A and B will be measured. The observations were = completed as scheduled, and no problems were reported. 1.5 Completed FOS/RD 6912 (FOS Cycle 6: Dark Monitoring) The Faint Object Spectrograph (Red Side) was used to make = measurements of the internal background. All exposures were internals, = and the spectrographic observations were made in the image mode. The = observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: There were no significant operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH